OpenStudy (lolo):

7th grade algebra? example:5x+10=45x thats just an example

OpenStudy (anonymous):

u wanna find x?

OpenStudy (anonymous):

So you want to get a tutor. Right?

OpenStudy (lolo):

yes

OpenStudy (anonymous):

And what type. Paid or free?

OpenStudy (lolo):

free

OpenStudy (anonymous):

Well, then you will have to post the problems here, and someone will reply.

OpenStudy (anonymous):

You can't expect to get help from the same person

OpenStudy (anonymous):

5x+10=45x =>10=45x-5x =>10=40x =>1/4=x

OpenStudy (lolo):

-3+Y=-5

OpenStudy (lolo):

i need help with intergers

OpenStudy (anonymous):

Please post them as separate questions. As I told you, one can't help you through out. There is a time factor

OpenStudy (lolo):

-10+4=N

OpenStudy (anonymous):

please post them as separate questions. As I told you, one can't help you through out. There is a time factor